OpenStudy (anonymous):

Is the expression x^3•x^3•x^3 equivalent to x^3•3•3?

jimthompson5910 (jim_thompson5910):

Do you mean "Is the expression \(\large x^3*x^3*x^3\) equivalent to \(\large x^{3*3*3}\) ?"

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

One of the many algebraic rules is that \[\Large x^a*x^b = x^{a+b}\] notice how the exponents 'a' and 'b' are added together. They are NOT multiplied You can extend the rule out to as many exponents as you want \[\Large x^a*x^b*x^c = x^{a+b+c}\] \[\Large x^a*x^b*x^c*x^d = x^{a+b+c+d}\] \[\Large x^a*x^b*x^c*x^d*x^e = x^{a+b+c+d+e}\] etc etc just make sure that the bases are all the same number

jimthompson5910 (jim_thompson5910):

just make sure that the bases are all the same number or algebraic expression (like x or something)

OpenStudy (anonymous):

Thank you very much!

jimthompson5910 (jim_thompson5910):

no problem